OpenStudy (anonymous):

I WILL MEDAL! How many solutions does the equation have? a + 5 = 1/5(5a + 25) A.one solution B.infinite solutions C.no solution

OpenStudy (anonymous):

a+5=(5a+25)/5 a-5=a+25 a-a=5+25 0=30 C - no solution is the correct answer

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

whenever you need help, feel free to tag me :)